# Super-Resolved Canopy Height Mapping from Sentinel-2 Time Series Using LiDAR HD Reference Data across Metropolitan France

### v1

@@ Abstract
Fine-scale forest monitoring is essential for understanding canopy structure and its dynamics, which are key indicators of carbon stocks, biodiversity, and forest health. Deep learning is particularly effective for this task, as it integrates spectral, temporal, and spatial signals that jointly reflect the canopy structure. To address this need, we introduce THREASURE-Net, a novel end-to-end framework for Tree Height Regression And Super-Resolution. The model is trained on Sentinel-2 time series using reference height metrics derived from LiDAR HD data at multiple spatial resolutions over Metropolitan France to produce annual height maps. We evaluate three model variants, producing tree-height predictions at 2.5 m, 5 m, and 10 m resolution. THREASURE-Net does not rely on any pretrained model nor on reference very high resolution optical imagery to train its super-resolution module; instead, it learns solely from LiDAR-derived height information. Our approach outperforms existing state-of-the-art methods based on Sentinel data and is competitive with methods based on very high resolution imagery. It can be deployed to generate high-precision annual canopy-height maps, achieving mean absolute errors of 2.62 m, 2.72 m, and 2.88 m at 2.5 m, 5 m, and 10 m resolution, respectively. These results highlight the potential of THREASURE-Net for scalable and cost-effective structural monitoring of temperate forests using only freely available satellite data. The source code for THREASURE-Net is available at: https://github.com/Global-Earth-Observation/threasure-net.
@@ 1 Introduction
Nowadays, forest monitoring is essential to capture the spatial heterogeneity in canopy structure and dynamics that drives carbon storage variability, species diversity, and early detection of forest health decline in the context of climate change (FAO, 2025). However, many forest processes – such as selective logging, canopy degradation, regeneration, or subtle structural changes – occur at fine spatial scales that cannot be captured by coarse imagery (Cally et al., 2025; Duncanson et al., 2025). Indeed, some vegetation monitoring tasks require finer spatial detail, especially in forests, where small-scale dynamics can indicate the early stages of irreversible change (Mitchell et al., 2017).
Canopy height prediction is a key component of forest monitoring, as it provides insights into stand structure, biomass, regeneration dynamics, and disturbance impacts (Lang et al., 2023). Although this task is not new (Simard et al., 2011), the methods and spatial resolutions involved have evolved substantially over time. While early approaches relied on coarse to medium resolution imagery combined with traditional machine learning models such as Random Forests (Morin et al., 2023; Ahmed et al., 2015; Wang et al., 2016), more recent works have shown that deep learning methods can predict canopy structure at much finer spatial resolutions, substantially improving height prediction accuracy (Fogel et al., 2025).
Moreover, forest dynamics are increasingly studied in their temporal context (Bugmann and Seidl, 2022), as analyzing changes over time is essential for understanding regrowth patterns, disturbance impacts, and long-term structural evolution. However, detecting small, subtle temporal changes remains particularly challenging, even though such early signals often mark the onset of potentially irreversible processes Liu et al. (2019). Therefore, enhancing canopy height prediction at fine spatial resolution and at frequent intervals remains an open challenge, especially in rapidly changing forest ecosystems where small-scale dynamics carry important ecological significance.
To monitor canopy evolution across large territories over time, researchers increasingly rely on satellite imagery as a practical source of information (Besic et al., 2025). While low- to medium-resolution data are freely available and provide broad spatial and temporal coverage, acquiring high-resolution reference data remains expensive and is often limited in availability (Fassnacht et al., 2023). Therefore, training models directly on very high–resolution (VHR) imagery at regional or national scales is often impractical and costly.
In this context, two central research questions remain open: how can prediction accuracy be further improved, and how can very high-resolution canopy height estimates be produced while keeping data acquisition costs low. To address these questions, we focus in this work on improving tree height prediction while simultaneously increasing the output resolution using only freely available remote sensing data.
In recent years, satellite data from missions such as Sentinel-1 and Sentinel-2 have been widely used to monitor and predict various forest structural and biophysical characteristics over time (Santoro et al., 2024; Mermoz et al., 2024). While Sentinel-2 offers dense time series capturing seasonal variations, most approaches focus on summarizing these observations to derive annual estimates of forest structure (Lang et al., 2023).
Nevertheless, manual annotation of forest stands at large scales to obtain reference data is impractical and often impossible. As a result, it has become common practice to use LiDAR (Light Detection and Ranging) data – whether airborne or satellite-based – to derive reference data. Among the most widely used sources are GEDI (Dubayah et al., 2020) – a spaceborne LiDAR mission providing discrete vertical profiles at 25-m resolution – and various airborne LiDAR campaigns.
@@
Today, airborne LiDAR data remains the most accurate and comprehensive source of information for forest structure monitoring and analysis (Coops et al., 2021). It provides detailed 3D point clouds whose spatial resolution – directly linked to point density– allows the precise characterization of forest vertical and horizontal structure. Therefore, various metrics derived from LiDAR data in raster shape at different resolutions often serve as reference data to train various machine learning models based on satellite imagery (Coops et al., 2021). In France, the LiDAR HD [^1] https://geoservices.ign.fr/lidarhd national campaign aims to cover the entire metropolitan territory with a minimum density of 10 pulses per m², which constitutes a key reference dataset for forest structural studies. However, despite its extensive coverage of French territory, each area is typically observed only once, which prevents any temporal analysis using LiDAR HD data alone. Hence, the high point density of the LiDAR HD data enables the extraction of forest metrics at fine spatial resolutions, providing valuable reference datasets for VHR tree height prediction.
Today, many researchers across different domains seek to bridge the gap between commonly available medium- and low-resolution data and the rare VHR data necessary for accurate modeling and analysis (Lepcha et al., 2023). Recent advances in deep learning and image processing have made it possible to enhance the spatial resolution of input data, opening new opportunities for both image and video analysis – from capturing finer details to enabling high-quality restoration (Wang et al., 2021).
This progress has also strongly influenced the field of remote sensing, where improving the spatial resolution of satellite imagery has become a central goal (Wang et al., 2022a; Vu-Duc et al., 2025). While medium resolution satellite imagery, such as Sentinel-2 data, is freely available and provides global Earth coverage with 5-days revisit period (Drusch et al., 2012), VHR images are typically restricted for commercial use or offer limited spatial and/or temporal coverage. Therefore, many recent algorithms aim to transfer these high resolution (HR) features to low resolution (LR) imagery to overcome this limitation (Wang et al., 2022b).
While most of the studies focus on a super-resolution (SR) task alone (Salgueiro Romero et al., 2022; Liu et al., 2025; Michel et al., 2025; Müller et al., 2020), some studies integrate SR with a real-life prediction task to exploit rich spatial features of recovered high resolution details. Most of those works focus on semantic segmentation since the fine-texture spatial details are particularly valuable for object delineation.
For instance, Lei et al. (2019) use a dataset of small RGB image patches depicting airplanes, ships, and oil tanks. Low-resolution inputs are generated by downsampling the original high-resolution patches by a factor of four, and a joint super-resolution and segmentation network is then trained. In Abadal et al. (2021), the authors propose a Dual Network for Super-Resolution and Semantic Segmentation with a feature-affinity loss. In their approach, Sentinel-2 imagery is first downsampled to 20 m to create LR–HR pairs, and the network is trained to perform both SR and land-cover semantic segmentation. Deng et al. (2025) introduce the ISASeg model for semantic segmentation of impervious surface areas at 1 m resolution, using Jilin-1 VHR imagery as the SR supervision dataset. Finally, Cao and Weng (2024) investigate a different prediction task – building height regression at 2.5 m – where Sentinel-2 images serve as input and Google Earth imagery provides the target data.
In this context, downsampling high-resolution optical images to create training pairs is a common practice, but it introduces unrealistic low-resolution inputs and removes important low-feature information. In contrast, obtaining genuine high-resolution datasets is often prohibitively costly, limiting the availability of accurate ground-truth data.
As shown, state-of-the-art methods focus mostly on joint Single Image Super Resolution (SISR) analysis since their prediction tasks itself do not benefit from multi-temporal data. However, for certain applications, in particular, for vegetation analysis, capturing phenological variation across time is crucial for accurate prediction (Pelletier et al., 2018).
In contrast, Multi-Image Super-Resolution (MISR) techniques aim to jointly exploit multiple observations of the same area to reconstruct a higher resolution image. These methods have demonstrated significant gains in domains such as video restoration or change detection, yet they are rarely explored in the context of Earth Observation time series.
While we have not found any works performing simultaneously MISR and an associated downstream task, some works performing MISR alone can be cited. These approaches can be broadly categorized into: (i) temporal-transformer models that learn a shared temporal representation before upsampling (Okabayashi et al., 2024), (ii) diffusion-based models that reconstruct high-resolution outputs by iteratively denoising samples conditioned on the low-resolution inputs (Luo et al., 2024), and (iii) recursive fusion models (Deudon et al., 2020), that progressively merge the low-resolution observations.
@@
Today several SISR models (Michel et al., 2025; Cresson, 2022) can be exploited to super-resolve Sentinel-2 satellite image time series (SITS). Consecutively, this new VHR SITS can then be directly used as an input for a prediction algorithm, the computational cost of processing a super-resolved series is drastically higher than the low resolution ones. Moreover, errors in image calibration and co-registration within a time series can introduce significant misalignments with the reference data, leading to substantial inconsistencies in the super-resolved input time series.
Motivated by these findings, we propose a novel end-to-end framework for forest canopy height regression from Sentinel-2 time series while simultaneously performing super-resolution. Our model, named THREASURE-Net (Tree Height REgression And SUper REsolution), allows us to generate tree height maps at 2.5, 5 and 10 m spatial resolutions. Our algorithm performance outperforms state-of-the-art methods based on Sentinel data and is competitive with methods based on very high resolution imagery, and can be exploited to produce high-precision annual height maps.
@@
    THREASURE-Net trained solely on Sentinel-2 data achieves an accuracy comparable to models trained on VHR satellite data, while also providing coherent fine-scale spatial details that are not visible in predictions at the native resolution.
@@
    The model is able to handle irregular Sentinel-2 time series acquired during the leaf-on season (from May to October of the same year) without explicitly requiring cloud-free observations.
@@ 2 Data
For model training and inference, we have prepared a training dataset spanning the entire area of metropolitan France, which corresponds to $80$ Sentinel-2 tiles following the Military Grid Reference System (MGRS). Each tile is divided into $1\times 1$km LiDAR HD (national LiDAR campaign that aims to cover the whole french territory with density 10 pts/m²) patches acquired over 2021-2024 years, resulting in $179945$ patches for training, $19133$ for validation, and $35639$ for testing. We spatially and temporally match each LiDAR HD patch with Sentinel-2 images time series. The spatial distribution of the reference LiDAR patches is shown in Figure 1.
To build Sentinel-2 SITS, we choose images acquired from May to October, on the same year as the reference LiDAR HD patch. We filter out images with cloud coverage and missing data superior to 50% and 10%, respectively.
Sentinel-2 image time series correspond to the Level 2A product, processed from L1C by MAJA  [^2] https://www.cesbio.cnrs.fr/maja/ algorithm developed by THEIA. We select 10 spectral bands (B02, B03, B04, B05, B06, B07, B08, B8A, B11, B12), Cloud Mask, as well as acquisition angles: Sun azimuth and zenith angles - $\theta_{S_{a}}$ and $\theta_{S_{z}}$ respectively, satellite mean view azimuth $\theta_{V_{a}}$ and zenith angles $\theta_{V_{z}}$. All bands and angles are resampled at 1010m. The acquisition angles are encoded using trigonometric transformations to respect angular periodicity:
@@
The reference data are derived from LiDAR HD 3D point clouds and transformed to height rasters in a fully-automatic way: point cloud values are projected onto a regular pixel grid of chosen spatial resolution, from which the $95^{th}$ percentile of elevation values is computed for every pixel. The LiDAR HD product is provided with point-wise classification into multiple land cover categories. Tree height values are only derived for pixels corresponding to vegetation higher than 1.5m, while the remaining pixels are set to zero. However, some tall vegetation crops, such as corn, may be classified as forest vegetation, introducing potential bias into model training, as crop vegetation does not follow the same structural or phenological characteristics as trees To solve this issue, we apply a crop mask derived from Land Parcel Identification System (Le Registre Parcellaire Graphique (RPG) [^3] https://artificialisation.developpement-durable.gouv.fr/bases-donnees/registre-parcellaire-graphique), the French national graphic crop register. All vegetation within crop parcels with height below 5m is considered non-permanent, and its height values are excluded from the annotations.

### v2

@@ Abstract
Fine-scale forest monitoring is essential for understanding canopy structure and its dynamics, which are key indicators of carbon stocks, biodiversity, and forest health. Deep learning is particularly effective for this task, as it integrates spectral, temporal, and spatial signals that jointly reflect the canopy structure. To address this need, we introduce THREASURE-Net, a novel end-to-end framework for Tree Height Regression And Super-Resolution. The model is trained on Sentinel-2 time series using reference height metrics derived from LiDAR HD data at multiple spatial resolutions over Metropolitan France to produce annual height maps. We evaluate three model variants, producing tree-height predictions at 2.5 m, 5 m, and 10 m resolution. THREASURE-Net does not rely on any pretrained model, nor on reference very high resolution optical imagery to train its super-resolution (SR) module; instead, it learns solely from LiDAR-derived height information. Our approach outperforms existing state-of-the-art methods based on Sentinel data and is competitive with methods based on very high resolution imagery. It can be deployed to generate high-precision annual canopy-height maps, achieving mean absolute errors of 2.63 m, 2.70 m, and 2.89 m at 2.5 m, 5 m, and 10 m resolution, respectively. These results highlight the potential of THREASURE-Net for scalable and cost-effective structural monitoring of temperate forests using only freely available satellite data. The source code for THREASURE-Net is available at: https://github.com/Global-Earth-Observation/threasure-net.
@@ 1 Introduction
Forest monitoring has always been essential for the sustainable functioning of ecosystems, managing the balance between harvesting for industry, carbon storage, and species diversity  (FAO, 2025). Canopy height mapping is a key component of forest monitoring, as it provides insights into stand structure, biomass, regeneration dynamics, and disturbance impacts (Lang et al., 2023). Since many forest processes – such as selective logging, canopy degradation, regeneration, or subtle structural changes – occur at fine spatial scales, methods and spatial resolutions have evolved substantially over time to improve the accuracy and spatial resolution of canopy height predictions.
@@
To monitor canopy evolution across large territories over time, researchers increasingly rely on satellite imagery as a practical source of information (Besic et al., 2025). Initially, a statistical approach was used to combine satellite imagery with in situ measurements (Simard et al., 2011; Avitabile et al., 2016; Bouvet et al., 2018). The in situ or field measurements can be national inventories carried out by the national forest agency at regular intervals (every 5 years in France) or isolated campaigns. Nevertheless, manual annotation of forest stands at large scales to obtain reference data is impractical and costly. As a result, it has become common practice to use LiDAR (Light Detection and Ranging) data – whether airborne or satellite-based – to derive reference data. Among the most widely used sources are GEDI (Dubayah et al., 2020) a spaceborne LiDAR mission providing discrete vertical profiles at 25-m resolution, and various airborne LiDAR campaigns. Today, airborne LiDAR data remains the most accurate and comprehensive source of information for forest structure monitoring and analysis (Coops et al., 2021). It provides detailed 3D point clouds whose spatial resolution – directly linked to point density – allows the precise characterization of forest vertical and horizontal structure. Many European countries have launched national LiDAR monitoring campaigns and provide Airborne laser scanning (ALS) data free of charge (Kakoulaki et al., 2021). In France, the LiDAR HD [^1] https://geoservices.ign.fr/lidarhd national ALS campaign aims to cover the entire metropolitan territory with a minimum density of 10 pulses per m², which constitutes a key reference dataset for forest structural studies. However, despite its extensive coverage of French territory, each area is typically observed only once, which prevents any temporal analysis using LiDAR HD data alone.
To overcome the lack of temporal information and spatial discontinuity, these LiDAR datasets are used in conjunction with satellite image series to produce temporally continuous maps. The methods and spatial resolutions involved have evolved substantially over time. While early approaches relied on coarse to medium resolution imagery combined with traditional machine learning models such as Random Forests (Morin et al., 2023; Ahmed et al., 2015; Wang et al., 2016), more recent works have shown that deep learning methods using high resolution (HR) open access satellite data from missions such as Sentinel-1 and Sentinel-2, at 10-20 m of resolution, can predict canopy structure with greater accuracy (Lang et al., 2023; Schwartz et al., 2023; Becker et al., 2023). These methods can produce annual maps, thereby increasing the frequency of forest monitoring. To overcome the limits of the 10 m spatial resolution, recent studies use very high-resolution (VHR) commercial imagery such as PlanetScope (Liu et al., 2023), Maxar (Tolan et al., 2024), or SPOT-6/7 (Fogel et al., 2025). However, acquiring VHR imagery remains expensive; therefore, training models at national scales is often impractical and costly.
In this context, two central research questions remain open: how can prediction accuracy be further improved, and how can very high-resolution canopy height estimates be produced while keeping data acquisition costs low. To address these questions, we focus in this work on improving tree height prediction while simultaneously increasing the output resolution using only freely available remote sensing data and super-resolution deep learning methods.
Super-resolution algorithms aim to transfer high resolution (HR) features to low resolution (LR) imagery  (Wang et al., 2022). While most studies focus solely on the super-resolution (SR) task (Salgueiro Romero et al., 2022; Liu et al., 2025; Michel et al., 2025; Müller et al., 2020), some studies combine SR with a real-life prediction task (such as building height estimation or object segmentation) to leverage the additional spatial information from higher-resolution images (Lei et al., 2019; Abadal et al., 2021; Cao and Weng, 2024). In these approaches, the SR component is typically supervised using either VHR images or artificially downsampled data to create LR–HR training pairs, while a separate dataset is used to supervise the downstream task.
@@
These methods focus on joint Single Image Super Resolution (SISR) analysis since their prediction tasks do not benefit from multi-temporal data. Today, several SISR models (Michel et al., 2025; Cresson, 2022) can be used to enhance Sentinel-2 satellite image time series (SITS). Therefore, the resulting VHR time series can then be directly used as input for prediction algorithms. However, processing super-resolved image series requires significantly more computational resources compared to using the original low-resolution data. Moreover, errors in image calibration and co-registration within a time series can introduce significant misalignments with the reference data, leading to substantial inconsistencies in the super-resolved input time series, potentially reducing the reliability of the prediction results.
In contrast, Multi-Image Super-Resolution (MISR) techniques aim to jointly exploit multiple low resolution observations of the same area to reconstruct a higher resolution image. While these methods have demonstrated significant gains in domains such as video restoration or change detection (Wang et al., 2021), they are rarely explored in the context of Earth Observation. Furthermore, they are better suited to capture phenological variation over time, which is crucial for accurate vegetation analysis (Pelletier et al., 2018).
While we have not found any works performing MISR simultaneously with a downstream prediction task, some works focusing solely on MISR can be cited. These approaches can be generally categorized into three categories: (i) temporal-transformer models that learn shared information across time before increasing the image resolution  (Okabayashi et al., 2024), (ii) diffusion-based models that reconstruct high-resolution images by iteratively denoising samples while being guided by the low-resolution inputs (Luo et al., 2024), and (iii) recursive fusion models (Deudon et al., 2020), that progressively combine multiple low-resolution observations to generate a higher-resolution output.
Motivated by these findings, we propose a novel end-to-end framework that simultaneously performs super-resolution and forest canopy height estimation using Sentinel-2 image time series. Our model, named THREASURE-Net (Tree Height REgression And SUper REsolution), allows us to generate tree height maps at 2.5, 5 and 10 m spatial resolutions. We apply our method to metropolitan France since forest monitoring in France still relies largely on field measurements and photointerpretation of high-resolution imagery, compared with Scandinavian countries, where LiDAR and satellite imagery are already integrated into national forest inventory systems. The objective of this study is therefore to provide new tools that can complement existing practices in France. To assess their relevance, we compare our approach with several state-of-the-art algorithms commonly used for forest variable estimation from satellite data.
@@
    THREASURE-Net, trained solely on Sentinel-2 data, achieves an accuracy comparable to models trained on VHR satellite data, while also providing coherent fine-scale spatial details that are not visible in predictions at the native resolution.
@@
    The model is able to handle irregular Sentinel-2 time series without explicitly requiring cloud-free observations.
  • 5.
    Our approach simultaneously predicts both canopy height and a tree/non-tree classification mask, whereas existing methods restrict height prediction to pre-defined forest areas using an external mask.
@@ 2 Data
2.1 Study area
Our area of study is mainland France (including Corsica), which spans 543,940 km². French forests cover 17.5 million hectares, accounting for nearly one-third of the territory. This makes them the fourth largest in Europe by surface area, after Sweden, Finland, and Spain. Three-quarters of these forests are privately owned and are mainly concentrated in the southern and eastern parts of the country. France’s geographical diversity is notable, featuring five distinct climates (oceanic, transitional oceanic, semi-continental, Mediterranean, and mountain) and $190$ tree species, as well as various silvicultural techniques. This diversity makes it a prime example of temperate European forests. The National Forest Inventory has divided the metropolitan territory into $91$ sections called sylvoécorégions (SER) – forest ecoregions – taking into account the bio-geographical factors that determine forest production and the distribution of major forest habitats (e.g., altitude, climate, and soil characteristics), see Figure 10.
Table 1: Overview of datasets used in this study.
Data Time span Resolution Description Source Usage
Sentinel-2 SITS 2018–2024 10 m Multispectral satellite imagery providing dense time series used to capture spectral and temporal information related to canopy structure. Copernicus Training / Evaluation
LiDAR HD 2021–2024 10 pts/m2 Airborne LiDAR dataset providing high-resolution canopy height metrics used as reference data for model training and evaluation. IGN Training / Evaluation
Land Parcel Identification System (RPG) 2024 Agricultural parcel dataset used to identify and mask non-forest areas within the study region. RPG – Data.gouv Masking
FORMS-T Schwartz et al. (2025) 2018–2023 10 m Canopy height prediction maps from FORMS-T algorithm for different years, used to compare with our predictions for independent ALS and disturbance detection. FORMS-T Comparison
Open-Canopy Fogel et al. (2025) 2018–2023 1.5 m Canopy height prediction maps from Open-Canopy algorithm for different years, used to compare with our predictions for independent ALS and disturbance detection. FORMSpoT Comparison
Global Canopy Height (GCH) Lang et al. (2023) 2020 10 m Canopy height prediction maps from GCH model used as reference for 2020 to compare with our predictions. GCH Model Comparison
Carcans-Hourtins 10/2020 31 pts/m2 ALS-derived tree height map. Coastal plain, pine forest with a few broadleaf trees in the understory. ONF Evaluation
Déodatie 04/2018 13 pts/m2 ALS-derived tree height map. Hilly terrain, mixed forest: beech, oak, hornbeam, spruce, and pine. ONF Evaluation
Lajoux-Lafresse 06/2019 31 pts/m2 ALS-derived tree height map. Mountainous terrain, spruce, pine, and beech. ONF Evaluation
Mouterhouse 02/2019 17 pts/m2 ALS-derived tree height map. Flat plain with small hills, beech, oak, hornbeam, and pine. ONF Evaluation
2.2 Training dataset
To train THREASURE-Net, a dataset composed of Sentinel-2 SITS and corresponding LiDAR HD patches was generated.
The French territory is covered by $80$ Sentinel-2 tiles following the Military Grid Reference System (MGRS). To build Sentinel-2 SITS, we choose images acquired from May to October, on the same year as the reference LiDAR HD patch. We filter out images with cloud coverage and missing data superior to 50% and 10%, respectively.
Sentinel-2 image time series correspond to the Level 2A product, processed from L1C by MAJA  [^2] https://www.cesbio.cnrs.fr/maja/ algorithm developed by THEIA, the French land data center for Earth observation. We select 10 spectral bands (B02, B03, B04, B05, B06, B07, B08, B8A, B11, B12), Cloud Mask, as well as acquisition angles: Sun azimuth and zenith angles - $\theta_{S_{a}}$ and $\theta_{S_{z}}$ respectively, satellite mean view azimuth $\theta_{V_{a}}$ and zenith angles $\theta_{V_{z}}$. All bands and angles are resampled at 10 m. The acquisition angles are encoded using trigonometric transformations to respect angular periodicity:
@@ 2.2 Training dataset
The reference data are derived from the Institut Geographique National (IGN)’s LiDAR HD open dataset, with patches acquired between 2021 and 2024. This consists of a national airborne LiDAR campaign that aims to cover the entire French territory with a minimum density of 10 pts/m². Patches have already been acquired over 2021-2025 years, and the campaign is ongoing. The 3D point clouds are transformed into height rasters in a fully-automatic way: point cloud values are spatially projected onto a regular pixel grid with a chosen spatial resolution (10 m, 5 m or 2.5 m), from which the $95^{th}$ percentile of elevation values is computed for the ensemble of points falling within each pixel. The LiDAR HD product is provided with point-wise classification into multiple land cover categories. Tree height values are only derived for pixels corresponding to vegetation higher than 1.5 m, while the remaining pixels are set to zero. However, tall crop vegetation, such as corn, may be incorrectly classified as forest vegetation. This introduces potential bias into model training, as crop vegetation does not exhibit the same structural or phenological characteristics as trees. To solve this issue, we apply a crop mask derived from Land Parcel Identification System (Le Registre Parcellaire Graphique (RPG)), the French national graphic crop register. All vegetation within crop parcels with height below 5 m is considered non-permanent, and its height values are excluded from the annotations.
To ensure a spatially robust and statistically representative dataset, we implemented a stratified sampling strategy on the LiDAR HD tiles, taking into account forest coverage and canopy height variance in the SER regions. First, a subset of 15% of LiDAR HD tiles was assigned to the test dataset, providing an independent evaluation dataset. To reduce spatial autocorrelation and prevent information leakage, all LiDAR tiles adjacent to the test tiles, including diagonal neighbors, were designated as buffer zones and excluded from the training and validation sets. The remaining tiles were randomly split into training (90%) and validation (10%) subsets. Within this stratified dataset, samples were randomly selected to obtain $179945$ patches for training, $19133$ for validation, and $35639$ for testing. This approach ensures the spatial independence of the test dataset while preserving regional representativeness in model development.
2.3 Ancillary data
To compare our canopy height estimates with independent reference ALS data, we use four height maps with 1 m resolution, derived from various high-density LiDAR campaigns conducted by the Office National des Forêts (ONF).
To compare our algorithm with the concurrent approaches, we equally use canopy height prediction maps provided by the authors FORMS-T (Schwartz et al., 2025) at 10 m resolution, Open-Canopy (Fogel et al., 2025) at 1.5 m resolution for different years, and the Global Canopy Height (GCH) Model (Lang et al., 2023) at 10 m resolution for 2020 year.
FORMS-T (Schwartz et al., 2025) uses yearly median composites of Sentinel-1 and Sentinel-2 imagery and GEDI LiDAR data as reference to predict canopy structure at 10 m resolution over the French metropolitan territory. In contrast, Open-Canopy (Fogel et al., 2025) relies on a single VHR SPOT-6/7 image combined with LiDAR HD to map fine-scale canopy height at 1.5 m, also over the French metropolitan territory. Finally, the GCH model provides global 10 m canopy height estimates using GEDI LiDAR as a reference and a single-date Sentinel-2 image, combined with coordinate encoding to identify geographical zones as an input.
A summary of all datasets used in this study is provided in Table 1.